X-ray photons are produced by electrons accelerated through a potential difference of 63 kV. Determine the wavelength of such a photon in nm. Your answer should be a number with four decimal places, do not include unit.

1nm = 10-9 m
Answer #1

V = 63kv 63000v

Since E = hc/lemda = eV

=) Lemda = hc/eV = 1240/63000 nm = 0.01968nm

=) Wavelength lemda = 0.0196nm